OpenStudy (anonymous):

(4x+7) - (4x+3) solved out ?

OpenStudy (anonymous):

4x+4x=4x 7+3=10 4x+10

OpenStudy (anonymous):

what do you think

OpenStudy (anonymous):

it's just 4

OpenStudy (anonymous):

thanks and yeah its 4

OpenStudy (anonymous):

and the X?

OpenStudy (anonymous):

i think is 4x+10

OpenStudy (anonymous):

its polynomials for subtracting its only 4

OpenStudy (anonymous):

no x 4x+7-(4x)-(3) 4x-4x=0 7-3=4